OpenStudy (anonymous):

Suppose 3s represents an even integer. What polynomial represents the product of 3s, the even integer that comes just before 3s, and the even integer that comes just after 3s? A. 27s3 + 12s B. 27s3 - 3s C. 27s3 - 12s D. 3s3 - 12s

OpenStudy (midhun.madhu1987):

 so (3s-2)(3s)(3s+2) = (3s-2)(3s+2)(3s) = ((3s)^2 - (2)^2)(3s) = (9s^2 - 4)(3s) = ?